P2.13 (similar to) -Question Help * An investor buys 200 shares of stock selling at $90 per share using a margin of 62% The stock pays annual dividends of $2 00 per share A margin loan can be obtained at an annual interest cost of 8.4%. Determine what return on invested capital the investor will realize if the price of the stock increases to $107 within six months. What is the annualized rate of return on this transaction? If the price of the stock increases to $107 within six months, the sik-month return on this transaction is% (Round to two decimal places.)
Answer #1

Loan Amount = 90*200*(1-Margin) = 18000*(1-62%) = 6840
Margin Amount = 18000*62% = 11,160

Interest Paid half yearly= Loan * Annual interest/2 = 6840*8.4%/2 = 574.56/2 = 287.28

Rise in price of stock = 107*200 = 21400
Dividend half yearly = 2*200/2 = 200
SIx month Return % = ( New Stock Value - Initial Stock Value - Interest + Dividend )/Margin amount = (21400-18000-287.28 +200)/11160 = 29.68%

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well